C1 220 nF C2 220 nF LM324 R1 12 k2 VA R2 12 k2 C3 440 nF Fig. 1-Twin T Notch Filter fo = 1/21*R1*c1

Using nodal analysis, calculate the transfer function of notch filter. Use the transfer function to calculate the expected gain (using the actual resistor and capacitor values that were measured for your experiment-See below). Hint: You need to solve the circuit with nodal analysis, using the impedance of a capacitor as - j / (2 * pi * C). The amplifier at the end is just a unity gain amplifier with a gain of 1, so it won't enter into the calculation.

R1=11.91kΩ, R2=11.95kΩ, R3=12.14kΩ, R4=11.94kΩ

C1=192nF, C2=192nF, C3=190nF, C4=195nF
Answer #1

oo figftly convoord concert all to reaitance ar gronf »x, :-) 274,6 to . 2TIRICI 2x3.14 X E XIC xp20 Xic Tag : 19068443 76kuat node 4, 3-1, of as of Py = 0 DAGER - 16-11 4-1 - Va t 1 - + - = 0 -jøk i Pare - jók ok IAI - 1 ta to I GK PRE-JEK 6K PRE-j>> VC140°0384 - 3003836) = 1; CO-038 +j0.1994 +0.0384 -30°03824) lo (0.2697 10.346 +0.1153 + 10.0385 ) ? Vo (10384 -0.03836)Nos [v; 6-10.178 - 101905) + 16 COV8 29-30.415))*(+30:5002 + [v; CO-0679 +30:1509) +V0C0-8559 + j0-384))x (0500?0.5002) # Vo